# Surface-Aware Distilled 3D Semantic Features

### v1

@@ 5.3 Ablations
We demonstrate the effectiveness of our hyperspherical embedding by replacing our contrastive loss $\mathcal{L}_{c}$ (Eq. 2) with three different options inspired by related work (see Appendix 8.2). First, the Relative Geodesic Loss (RGL[41] optimizes relative distances in a Euclidean embedding. Similarly, the Naive Geodesic Loss (NGL) minimizes absolute distances. Finally, the Geometrical Similarity Loss (GSL[14] enforces similarity of feature and surface distances in a local neighborhood. We remove feature and geodesic normalization wherever absolute magnitude needs to be learned. In Tbl. 2 (top), we observe that, except for correspondence accuracy for TOSCA, our contrastive loss $\mathcal{L}_{c}$ outperforms either of the alternatives in the correspondence matching task.
@@
In Tbl. 2 (bottom), we individually assess our two losses. We see that the performance with only the reconstruction loss $\mathcal{L}_{r}$ is close to Diff3F. This indicates that the gain in performance does not originate predominantly from a smaller embedding space or from access to training data. Similarly, the contrastive loss $\mathcal{L}_{c}$ alone results in an accuracy drop compared to our full model. This justifies our autoencoder approach with both losses playing an import role.
Number of Anchors
We train our method for different anchor counts $A$ with a constant two-hour training budget per model. In Fig. 8, we observe a low sensitivity to the anchor count. Due to repeated random sampling over the course of training, even $A=1$ outperforms the correspondence error of Diff3F. Ultimately, we opt for $A=100$ in all our experiments, as it balances computation cost and matching performance well. Note that $A$ is not used during inference.
@@
Table 2: Ablation of our method. Above the bar: Ablation of alternative losses inspired by related work [14, 41] compared to the unmodified Diff3F features. Below the bar: Our full method compared to its reduced variant omitting losses $\mathcal{L}_{c}$ or $\mathcal{L}_{r}$. The per-column best results are bold and the second-to-best are underlined.
Figure 8: Ablation of anchor count $A$ on SHREC’19. Our method improves upon Diff3F already from $A=1$.
6 Applications
@@
We present additional downstream tasks that benefit from surface-aware features learned in Sec. 5.1.
6.1 Instance-based Part Segmentation
@@
Following prior work [24], we segment a target shape by clustering features around centroids from K-means clustering of source-shape features. In Fig. 9, we demonstrate a transfer from a big cat to a human and see that unlike the Diff3F features, our surface-aware features disambiguate the limbs. In Fig. 1, we repeat this with centroids obtained jointly from all SHREC’19 and TOSCA samples. In contrast to Diff3F (see Appendix 10.2), our method successfully matches features across diverse shapes.
@@
Figure 9: Ten k-means cluster centers from the big cat were used to segment the human. Our method successfully separates all limbs.
6.2 Pose Alignment
Our surface-aware features are also useful for the pose alignment of a kinematic model to another 3D shape. To this end, we establish point correspondences between shape pairs as in Sec. 5.1 and optimize the kinematic pose parameters to minimize point-to-point distances (see Appendix 9).
In Fig. 10, we align SMPLeval to SHREC’19, and SMALeval to DeformingThings4D [45] animals. Benefiting from the robust instance separation, our method produces poses closer to the targets for both dense and sparse correspondences. See our video for a 3D shape animation obtained by aligning to a target shape sequence.
@@ 6.2 Pose Alignment
Figure 10: Pose alignment of a source shape (color) into the pose of a target (gray). The boxes highlight challenging areas handled well by our method. Top two rows densely fit to all target vertices, the bottom three only fit to 5 % of the target vertices (highlighted).
7 Discussion
@@
Our method inherits limitations connected to the extraction of the base features. Specifically, the extraction of Diff3F features [24] takes several minutes per mesh, and its vision model is sensitive to rendering artifacts. We expect that advances in rendering of point representations increase the applicability across representations [43]. Furthermore, objects that are isotropic both geometrically and semantically do not enable our method to establish a consistent partitioning of instanced components. Hence, while our embedding separates human legs following the body’s notion of front and rear, it cannot do so for table legs. However, this is not an issue for applications such as shape morphing [74].
Beyond 3D alignment, our methodology could inspire 3D-to-2D pose estimation [42, 59], articulated 3D reconstruction [86, 77], automated rigging [84] or 2D-to-3D uplifting [61, 47], where our features could support more view-consistent representations. Finally, an interesting topic for future research is the development of foundational features using massive datasets, such as Objaverse [18].
@@ Conclusion
We have introduced novel surface-aware features for 3D shape matching that disambiguate intraclass instances among semantic features derived from pre-trained 2D vision models. Our descriptors have proven effective in distinguishing instances of the same semantic class. They are capable of generalization even when trained on a limited number of 3D shapes. Furthermore, our contrastive loss facilitates easy integration in future methods without data labeling. Consequently, we perceive our method as a promising building block in the research landscape, aiming to exploit pre-trained 2D models for 3D tasks.
Supplementary Material
8 Additional implementation details
We provide additional details to reproduce our experiments. We will make the code publicly available upon publication.
8.1 Diff3F
We use the authors’ code to compute the Diff3F features [24] for our base features and as a baseline method for comparison.
Since we observed that the camera poses used for sampling in the TOSCA dataset are biased towards a specific up-direction, we modified the code to flip the coordinate system for only this dataset. Doing so yields $\approx 10$% increase in correspondence accuracy in TOSCA for both our method and the Diff3F baseline, when compared to the numbers reported in the Diff3F paper [24]. An alternative solution could be a modification of the camera sampling algorithm itself.
8.2 Ablations
Utilizing the geodesic distance as a supervision signal for embeddings has been explored in related works  [41, 14]. However, in our ablations, we show that we achieve superior results with our formulation. In this section, we discuss the key differences.
@@ 8.2 Ablations
Based on two points, $v_{p}$ and $v_{q}$, the Relative Geodesic Loss [41] minimizes the difference between the geodesic distance $d^{S}$ and the Euclidean embedding distance $d^{E}$ of those two points, relative to the surface distance:
$L_{RGL}=\sum_{i}\sum_{(p,q)\in S_{i}}\frac{|d_{i}^{E}(v_{p},v_{q})-d^{S}(v_{p},v_{q})|^{2}}{d^{S}(v_{p},v_{q})^{2}}.$ (4)
The normalization term is introduced to prioritize local distance preservation. We do not utilize this normalization term, as the base features struggle to disambiguate samples that are far away on the surface but close in feature space, such as “left hand” and “right hand”.
@@
NGL is discussed in [41] but not used for training, as the authors state that it might hamper local distance preservation. Indeed, in our ablations, it achieves worse results in terms of correspondence accuracy when compared to RGL. It is identical to RGL but it omits the normalization term:
$L_{NGL}=\sum_{i}\sum_{(p,q)\in S_{i}\in[n_{i}]^{2}}|d_{i}^{E}(v_{p},v_{q})-d^{S}(v_{p},v_{q})|^{2}$ (5)
While not actually utilized in their work, NGL is closest to our formulation. However, our choice of a hyperspherical rather than Euclidean embedding space in combination with our autoencoder setup achieves notably better results.
@@
A concurrent work proposes to maximize the local angular similarity between geodesic distances and Euclidean feature vectors for a set of neighbors for each point [14]. Crucially, the similarity is only maximized for a set of $k$ point pairs nearest in the embedding space. The loss is a cosine metric between a vector of geodesic distances $\mathbf{m}_{i}\in\mathbb{R}^{k}$ and a vector of Euclidean embedding distances $\mathbf{d}_{i}\in\mathbb{R}^{k}$:
$\mathcal{L}_{\text{GSL}}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{\mathbf{d}_{i}\cdot\mathbf{m}_{i}}{\|\mathbf{d}_{i}\|\|\mathbf{m}_{i}\|}\right).$ (6)
@@ Conclusion
Our method differs from recent works utilizing vision-based features for 3D shape matching in three main aspects: 1) we follow a global approach when enforcing distances in the embedding space; 2) our embedding space is hyperspherical and it only encodes angular information; 3) in context of the whole pipeline, we solely rely on intrinsic properties.
9 Pose Alignment
We establish correspondences based on $\phi$, such that each point $\mathbf{x}_{i}^{\mathcal{S}}$ in the source shape is assigned a target point $\mathbf{x}_{i}^{\mathcal{T}}$ in the target shape (see Sec. 3). We then align the source to the target by minimizing the following L1 loss:
$\mathcal{L}_{point}=\frac{1}{N}\sum_{i}^{N}||\mathbf{x}_{i}^{\mathcal{S}}-\mathbf{x}_{i}^{\mathcal{T}}||_{1}.$ (7)
For the first half of the optimization, we only optimize the root rotation $\mathbf{R}$, root translation $\mathbf{t}$, and scale $s$, which roughly aligns the meshes. In the second half, we additionally optimize the rotation $\mathbf{R}_{b}$ of each kinematic bone $b$. The parameters are optimized based on a gradient-descent for $4000$ iterations, which takes approximately 30 seconds for a static pose.
Furthermore, we found it beneficial to use an as-rigid-as-possible regularization [71], which penalizes the deviation between the initial edge lengths of the mesh $\delta^{init}_{e}$ and the current edge length $\delta_{e}$ for each edge $e$:
$\mathcal{L}_{arap}=\frac{1}{E}\sum_{e}^{E}|\delta_{e}^{init}-\delta_{e}|.$ (8)
@@ 9 Pose Alignment
$\mathcal{L}_{smooth}=\frac{1}{N(T-1)}\sum_{t}^{T-1}\sum_{i}^{N}||\mathbf{x}_{i,t}^{\mathcal{S}}-\mathbf{x}_{i,t+1}^{\mathcal{S}}||_{2}^{2}.$ (9)
@@
10 Additional results
10.1 Qualitative Results on SHREC’20
Fig. 11 presents additional results for the SHREC’20 dataset. As the dataset only provides $\approx 50$ correspondences for each shape pair, we do display dense ground-truth labels. However, we find that our features generally produce smoother correspondences (e.g., bottom left) and a better separation of legs (e.g., second to last row on the right).
Figure 11: Qualitative results on SHREC’20. Our features result on average in more accurate smooth mapping between different limbs.
10.2 Clustering
Same as in Sec. 6.1, we obtain the centroids from all SHREC’19 and TOSCA samples’ features, but this time with the Diff3F. We use those centroids to segment a set of shapes. The results are shown in Fig. 12. While animal limbs and and bodies are still assigned to the same cluster, limbs are not disambiguated and they are not assigned to the same clusters between humanoids and animals. Notably, this is not an issue with our surface-aware features in Fig. 1.
Figure 12: Reproduction of the clustering experiment shown in Fig. 1 but utilizing Diff3F features instead of our features. Notice the inconsistency of labeling across different shape categories as well as lack of separability between instanced components such as individual limbs.

### v2

@@ 5.5. Ablations
We demonstrate the effectiveness of our hyperspherical embedding by replacing our contrastive loss $\mathcal{L}_{c}$ (Eq. 2) with three different options inspired by related work (see Appendix A.2). First, the Relative Geodesic Loss (RGL(Jiang et al., 2023) optimizes relative distances in a Euclidean embedding. Similarly, the Naive Geodesic Loss (NGL) minimizes absolute distances. Finally, the Geometrical Similarity Loss (GSL(Chen et al., 2025) enforces similarity of feature and surface distances in a local neighborhood. We remove feature and geodesic normalization wherever absolute magnitude needs to be learned. In Tbl. 2 (top), we observe that, except for correspondence accuracy for TOSCA, our contrastive loss $\mathcal{L}_{c}$ outperforms all of the alternatives in the correspondence matching task.
@@
In Tbl. 2 (bottom), we individually assess our two losses. We see that the performance with only the reconstruction loss $\mathcal{L}_{r}$ is close to Diff3F. This indicates that the gain in performance does not originate predominantly from a smaller embedding space or from access to training data. Similarly, the contrastive loss $\mathcal{L}_{c}$ alone results in an accuracy drop compared to our full model. This justifies our autoencoder approach with both losses playing an import role. Ablations on the number of anchors can be found inAppendix A.2.
Table 2. Ablation on our method. Above the bar: Ablation on alternative losses inspired by related work (Chen et al., 2025; Jiang et al., 2023) compared to the unmodified Diff3F features. Below the bar: Our full method compared to its reduced variant omitting losses $\mathcal{L}_{c}$ or $\mathcal{L}_{r}$. The per-column best results are bold and the second-to-best results are underlined.
@@
6. Applications
@@
We present additional downstream tasks that benefit from our surface-aware features learned in Sec. 5.
6.1. Instance-based Part Segmentation
@@
Following the prior work (Dutt et al., 2024), we segment a target shape by clustering features around centroids from K-means clustering of source-shape features. In the top row of Fig. 9, we demonstrate a transfer from a big cat to a human and see that, unlike the Diff3F features, our surface-aware features disambiguate the limbs. In the bottom two rows, we repeat this experiment with a shared encoder trained on human, animals, and a subset of ShapeNet (see Appendix D.4) where a true mapping cannot be defined but our method finds reasonable analogies between the classes.
In Fig. 11, we repeat this with centroids obtained jointly from all SMPLeval and TOSCA samples. In contrast to Diff3F, our method successfully matches features across diverse shapes, which demonstrates our embedding’s capability of many-to-many shape matching without any additional pairwise optimization. Finally, we show similar results for chairs and airplanes in Fig. 12.
@@
Figure 9. In the top row, 10 k-means cluster centers from the big cat were used to segment the human. In the bottom two rows, 8 k-means cluster centers from the animals were used to segment the chairs and airplanes with a shared encoder. Unlike Diff3F, our method successfully separates all limbs for a plausible mapping from animal limbs to human limbs, chair legs, or airplane wings.
@@
Figure 10. Pose alignment of a source shape (color) into the pose of a target (gray). The boxes highlight challenging areas handled well by our method. For humans, we densely fit all the target vertices, while for animals, we only fit 55 % of the vertices as highlighted.
6.2. Pose Alignment
Our surface-aware features are also useful for pose alignment of a kinematic model to another 3D shape. To this end, we establish point correspondences between shape pairs as in Sec. 5.4 and optimize the kinematic pose parameters to minimize point-to-point distances (see Appendix C).
In Fig. 10, we align SMPLeval to SHREC’19, and SMALeval to DeformingThings4D (Li et al., 2021) animals. Benefiting from the robust instance separation, our method produces poses closer to the targets for both dense and sparse correspondences. See our video for a 3D shape animation obtained by aligning to a target shape sequence.
Figure 11. Results when clustering features across all samples in SMPLeval and TOSCA. Our method implicitly aligns semantically-related regions (shown as the same colors) across diverse 3D shapes in a self-supervised manner (the top row). Diff3F produces inconsistent labeling across different shape categories as well as lack of separability between instanced components such as individual limbs (the bottom row).
Figure 12. Results when clustering the surface-aware features across chairs (top row) and airplanes (bottom row) from ShapeNet (Chang et al., 2015). Note, that while we use a single shared encoder for all humanoid and animal shapes, we train a separate encoder for each ShapeNet class due to the large domain gap.
6.3. Texturing
Since the base features are obtained from image models (see Sec. 3) and our pointwise encoder can process points sampled from a mesh as easily as pixels sampled from an image, we can establish correspondences between a 2D image and a 3D mesh. We demonstrate this by texturing 3D meshes from a masked target image and individually assign each vertex a color from the image pixel that maximizes the mutual feature similarity (Eq. 1) (see Fig. 13 and Appendix D.3). We observe that our features produce a more coherent mapping leading to a better preservation of the source appearance when compared to the Diff3F base features. In Fig. 14, we further show that textures can also be effectively transferred between two 3D shapes using a combination our surface-aware features with Functional Maps like in Sec. 5.4.
Figure 13. Texturing of 3D meshes from SMPLeval and TOSCA, based on a 2D image generated with ChatGPT. The appearance is transferred by establishing correspondence between the image features and 3D mesh features. In contrast to Diff3F, our surface-aware features represent the input image more faithfully.
Figure 14. Texturing of 3D meshes based on a source 3D mesh. The appearance is transferred based on correspondences established by combining our surface-aware features with Functional Maps. The data originate from SMPLeval, SMPLitex (Casas and Comino-Trinidad, 2023), and DeformingThings4D (Li et al., 2021).
7. Discussion
@@
Our method inherits limitations connected to the extraction of the base features. Specifically, the extraction of Diff3F features (Dutt et al., 2024) takes several minutes per mesh and its vision model is sensitive to rendering artifacts or upside-down mesh orientations. We expect that advances in rendering of point representations increase the applicability across representations (Kerbl et al., 2023). Furthermore, our method cannot establish a consistent partitioning for objects that are both geometrically and semantically isotropic (e.g., a round table). Hence, while our embedding separates human legs following the body’s notion of front and rear, it cannot do so for table legs. However, this is not an issue for applications such as shape morphing (Sun et al., 2024). Lastly, our method relies on consistency of geodesic distances between semantically distinct parts, and therefore it will be affected by geodesic shortcuts for partially blended parts in noisy 3D reconstructions (e.g., touching hands).
Beyond 3D alignment, our methodology could inspire 3D-to-2D pose estimation (Kanazawa et al., 2018; Peng et al., 2019), articulated 3D reconstruction (Yao et al., 2022; Uzolas et al., 2023), automated rigging (Xu et al., 2020) or 2D-to-3D uplifting (Poole et al., 2022; Liu et al., 2023), where our features could support more view-consistent representations. Finally, an interesting topic for future research is the development of foundational features using massive datasets, such as Objaverse (Deitke et al., 2023).
@@ Conclusion
We have introduced novel surface-aware features for 3D shape matching that disambiguate intra-class instances among semantic features derived from pre-trained 2D vision models. Our descriptors have proven effective in distinguishing instances of the same semantic class and they generalize even when trained on a limited number of 3D shapes. Furthermore, our contrastive loss facilitates easy integration in future unsupervised methods which reduces data labeling effort. Consequently, our method is a promising building block toward adapting pre-trained 2D models to 3D tasks.
Acknowledgements.
This work was partially supported by the Convergence AI Immersive Tech Lab at TU Delft.
Appendix A Additional implementation details
Here, we provide additional details to reproduce our experiments.
A.1. Diff3F
We use the authors’ code to compute the Diff3F features (Dutt et al., 2024) for our base features and as a baseline method for our comparisons.
Since we observed that the camera poses used for sampling in the TOSCA dataset are biased towards a specific up-direction, we modified the code to flip the coordinate system for only this dataset. Doing so yields $\approx 10$% increase in correspondence accuracy in TOSCA for both our method and the Diff3F baseline, when compared to the numbers reported in the Diff3F paper (Dutt et al., 2024). An alternative solution could be a modification of the camera sampling algorithm itself.
A.2. Ablations
Figure 15. Ablation on the anchor count AA on SHREC’19. In terms of the correspondence error, our method improves upon Diff3F already starting from $A=1$.
Number of Anchors
We train our method for different anchor counts $A$ with a constant two-hour training budget per model. In Fig. 15, we observe a low sensitivity to the anchor count. Due to repeated random sampling over the course of training, even $A=1$ outperforms the correspondence error of Diff3F. Ultimately, we opt for $A=100$ in all our experiments, as it balances computation cost and matching performance well. Note that neither the parameter $A$ nor any computation of geodesic distances in generally are used during inference.
Losses
Utilizing the geodesic distance as a supervision signal for embeddings has been explored in related works  (Jiang et al., 2023; Chen et al., 2025). However, in our ablations, we show that we achieve superior results with our formulation. In this section, we discuss the key differences.
@@ A.2. Ablations
Based on two points, $v_{p}$ and $v_{q}$, the Relative Geodesic Loss (Jiang et al., 2023) minimizes the difference between the geodesic distance $d^{S}$ and the Euclidean embedding distance $d^{E}$ of those two points, relative to the surface distance:
(4) $L_{RGL}=\sum_{i}\sum_{(p,q)\in S_{i}}\frac{|d_{i}^{E}(v_{p},v_{q})-d^{S}(v_{p},v_{q})|^{2}}{d^{S}(v_{p},v_{q})^{2}}.$
The normalization term is introduced to prioritize local distance preservation. We do not utilize this normalization term, because the base features struggle to disambiguate samples that are far away on the surface but close in feature space, such as “left hand” and “right hand”.
@@
NGL is discussed by Jiang et al. (2023) but not used for training, as the authors state that it might hamper local distance preservation. Indeed, in our ablations, it achieves worse results in terms of correspondence accuracy when compared to RGL. It is identical to RGL but it omits the normalization term:
(5) $L_{NGL}=\sum_{i}\sum_{(p,q)\in S_{i}\in[n_{i}]^{2}}|d_{i}^{E}(v_{p},v_{q})-d^{S}(v_{p},v_{q})|^{2}$
While not actually utilized in their work, the NGL formulation is the closest of the three to our own formulation. However, our choice of a hyperspherical rather than Euclidean embedding space in combination with our autoencoder setup achieves notably better results.
@@
A concurrent work proposes to maximize the local angular similarity between geodesic distances and Euclidean feature vectors for a set of neighbors for each point (Chen et al., 2025). Crucially, the similarity is only maximized for a set of $k$ point pairs nearest in the embedding space. The loss is a cosine metric between a vector of geodesic distances $\mathbf{m}_{i}\in\mathbb{R}^{k}$ and a vector of Euclidean embedding distances $\mathbf{d}_{i}\in\mathbb{R}^{k}$:
(6) $\mathcal{L}_{\text{GSL}}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{\mathbf{d}_{i}\cdot\mathbf{m}_{i}}{\|\mathbf{d}_{i}\|\|\mathbf{m}_{i}\|}\right).$
@@ Conclusion
Our method differs from recent and concurrent works utilizing vision-based features for 3D shape matching in three main aspects: 1) we follow a global approach when enforcing distances in the embedding space; 2) our embedding space is hyperspherical and it only encodes angular information; 3) in the context of the whole pipeline, we solely rely on intrinsic properties.
Appendix B One Shot Pose Transfer
We train an MLP to model the deformation between the paired input source mesh $\mathcal{M}_{train}^{src}$ and the output target mesh $\mathcal{M}_{train}^{tgt}$ obtained from SMPLevaland thus not used for training of our features. The paired training meshes share the same identity $\beta$ but they differ in poses $\theta$ such that
$\mathcal{M}_{train}^{src}$ $:=SMPL(\beta_{src},\theta_{src})$
$\mathcal{M}_{train}^{tgt}$ $:=SMPL(\beta_{src},\theta_{tgt}).$
We train an MLP $\mathcal{M}_{train}^{tgt}=\mathcal{J}(\mathcal{M}_{train}^{src},\mathcal{S_{M}^{\text{{src}}}})$ to produce the target pose mesh $\mathcal{M}_{train}^{tgt}$ conditioned on the source mesh $\mathcal{M}_{train}^{src}$ and its surface-aware features $\mathcal{S_{M}^{\text{{src}}}}$. These features are produced by our pre-trained general encoder $\mathcal{E}(.)$ from Sec. 5 without any further fine-tuning. Alternatively, other features are used for comparisons.
During training, $\mathcal{J}(.)$ learns a per-face residual function  (Muralikrishnan et al., 2024) to model a Neural Jacobian Field (Aigerman et al., 2022) while being supervised by MSE between the predicted re-posed mesh and the target $\mathcal{M}_{train}^{tgt}$.
Crucially, $\mathcal{J}(.)$ is defined on a per-face basis, which means that the input meshes used in test time do not need to have the same connectivity as the training mesh pair. In quantitative comparisons, we apply the learned mapping $\mathcal{J}(.)$ to input meshes of unseen identities $\mathcal{M}_{test}^{src}:=SMPL(\beta_{test},\theta_{src})$ and compare against their ground truths $\mathcal{M}_{test}^{tgt}:=SMPL(\beta_{test},\theta_{tgt})$. For other shape classes, we provide qualitative comparisons.
Appendix C Pose Alignment
We establish correspondences between two input shapes based on the feature cosine similarity $\phi(.)$ (Sec. 3), such that each point $\mathbf{x}_{i}^{\mathcal{S}}$ in the source shape is assigned a target point $\mathbf{x}_{i}^{\mathcal{T}}$ in the target shape. Next, we align the source to the target by minimizing the following L1 loss:
[…]
For the first half of the optimization steps, we only optimize the root rotation $\mathbf{R}$, the root translation $\mathbf{t}$, and the scale $s$, which roughly rigidly aligns the meshes. In the second half, we additionally optimize the rotation $\mathbf{R}_{b}$ of each kinematic bone $b$. The parameters are optimized based on a gradient-descent for $4000$ iterations, which takes approximately 30 seconds for a static pose.
Furthermore, we found it beneficial to use an as-rigid-as-possible regularization (Sorkine and Alexa, 2007), which penalizes the deviation between the initial edge lengths of the mesh $\delta^{init}_{e}$ and the current edge length $\delta_{e}$ for each edge $e$:
(8) $\mathcal{L}_{arap}=\frac{1}{E}\sum_{e}^{E}|\delta_{e}^{init}-\delta_{e}|.$
@@ Appendix C Pose Alignment
[…]
@@
Appendix D Additional results
D.1. Qualitative Results on SHREC’20
Fig. 16 presents additional results for the SHREC’20 dataset. As the dataset only provides $\approx 50$ correspondences for each shape pair, we display the predicted correspondences without dense ground-truth labels. However, we find that our features generally produce smoother correspondences (e.g., bottom left) and a better separation of legs (e.g., the second to last row on the right).
Figure 16. Qualitative results on SHREC’20. Our features result on average in a more accurate smooth mapping between different limbs.
D.2. Qualitative Results on ShapeNet
In Fig. 17, we show additional results for chairs and airplanes from ShapeNet (Chang et al., 2015). Since no dense ground truth labels are available, we show the predicted correspondences alone. We find that our surface-aware features achieve results better than the Diff3F baseline when separating the chair legs (see the top left row) and the airplane wings (see the top right row).
Figure 17. Qualitative results on ShapeNet (Chang et al., 2015). Our features result on average in a more accurate smooth mapping between chair legs and wings.
D.3. Texturing
We provide additional examples of 2D-to-3D texturing based on our own features in Fig. 18.
Figure 18. Texturing of 3D meshes from SMPLeval and TOSCA, based on a 2D image generated with ChatGPT. The appearance is transferred by establishing correspondence between the image features and 3D mesh features. Our method performs well even on incomplete meshes (notice the bear in the second row).
Table 3. Correspondence metrics measured for human and animal test shapes. The Specialized encoder was trained on a join set of humans and animals following our point-to-point correspondence experiment procedure in Sec. 5.4, and thus, the values match Tbl. 1 in the paper (see Ours). All-shape encoder was trained on a larger more generalized training set covering humans and animals but also chairs and airplanes.
Encoder SHREC’19 TOSCA SHREC’20
Specialized err ↓ 0.43±0.76 1.65±2.15 23.89±8.90
acc ↑ 28.78±9.30 29.35±14.53 73.97±26.47
All-shape err ↓ 0.56±1.03 1.62±2.08 4.37±9.47
acc ↑ 27.53±9.47 27.45±14.56 70.33±24.86
D.4. Using a Shared Encoder for All Shapes
In Sec. 5, we train a shared encoder for (SURREAL) and animals (SMAL) shapes as well as separate encoders for chair and airplanes (ShapeNet). Here, we follow the same procedure and train a new single shared encoder on a union of all these shapes and test it on human and animal shapes as in the paper Tbl. 1 to assess further generalization of our approach. The correspondence metrics in Tbl. 3 show that the all-shape encoder generally slightly under-performs the specialized encoders but it still improves upon the baselines (see Tbl. 1 in the paper). TOSCA is an exception, as the larger combined train set marginally reduces the error.
D.5. Ablating the Number of Training Shapes
In Tbl. 4, we explore how the number of training shapes affects the correspondence error with the same fixed validation set and training policy as in the main experiments. We vary the number of training samples while retaining a constant animal-to-human shape ratio. We find that just 2 training samples already decrease the error when compared to Diff3F. As expected, additional samples lead to further improvements for SHREC’19 and TOSCA. This trend is more subtle for SHREC’20.
Table 4. Training set size (columns) vs. correspondence error.
Diff3f 2 10 50 94
SHREC’19 1.69±1.44 1.32±1.22 1.31±1.31 0.48±0.85 0.43±0.76
TOSCA 4.51±5.48 3.75±3.50 2.60±2.74 1.84±2.47 1.65±2.15
SHREC’20 5.34±10.22 3.89±8.49 4.05±9.90 3.96±9.33 3.89±8.90